[Hallc_running] [New Logentry] RC update, Jul 8 2022

berdnik at jlab.org berdnik at jlab.org
Fri Jul 8 09:50:04 EDT 2022

Logentry Text:
1) MCC did attempt to pump up the current above 40uA at the beginning of the day shift yesterday, unfortunately, looks like the machine still facing issues to deliver a high current to hallC...
2) The beam was off for the second part of the day shift until approximately 20-00 due to Dominion power (off-site) transformer work and afterwords beam restoration.
3) Stable beam was restored at ~35uA over night and we finished Dummy target runs, 3/4 HMS trigger run and continuing production on LH2
4) At the morning MCC meeting we decided to stay on stable ~40uA beam over weekend in order to complete  Q2 = 6.00 W=3.19, x=0.39, SHMS center setting by Monday
5) For the moment we collected 22124/49500 pions
6) HallC magnets, target and etc. was "quiet" for the last 24 hours

We'll have a meeting today at 11am
Zoom - https://jlab-org.zoomgov.com/j/1607100885?pwd=ZFBsRTFsYm5Zb0laZ0tyZXNjcE...


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4012666
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220708/b8b61e1b/attachment.html>

More information about the Hallc_running mailing list